# Privacy for Quantum Annealing. Attack on Spin Reversal Transformations in the Case of Cryptanalysis

### v3

@@ 2.1 Quantum cloud services and privacy
The adversary is between the client and the quantum cloud service. All communication goes through the adversary, which has the ability to perform any operation on the received problem. This paper assumes that adversaries eavesdrop on intercepted communications and pass them forward. With the growing number of cloud service providers, the threat is increasing. Service providers are in high demand, and their availability is severely limited. This leads to a situation where a new provider offering competitive services can gain popularity rapidly. The internal structure of cloud services remains a mystery, and it is unknown if there is no eavesdropping between the client and the actual computer. Such a new, untrusted provider may be malicious, and communications between the user and the quantum computer may be intercepted.
@@ 5.1 Example parameterization
    $\begin{array}[]{ccc}h_{0}=\phi(1,0,0)=\begin{bmatrix}3\\ 5\\ 2\end{bmatrix},&h_{1}=\phi(0,1,0)=\begin{bmatrix}2\\ 5\\ 1\end{bmatrix},&h_{2}=\phi(0,0,1)=\begin{bmatrix}2\\ 4\\ 7\end{bmatrix}.\end{array}$
@@
    $\begin{array}[]{ccc}h_{0}-h_{\infty}=\begin{bmatrix}1\\ 1\\ 1\end{bmatrix},&h_{1}-h_{\infty}=\begin{bmatrix}0\\ 1\\ 0\end{bmatrix},&h_{2}-h_{\infty}=\begin{bmatrix}0\\ 0\\ 6\end{bmatrix}.\end{array}$
@@ 5.2 An illustrative attack on a scaled-down instance
[…]

### v4

@@ 2.1 Quantum cloud services and privacy
The adversary is between the client and the quantum cloud service. All communication goes through the adversary, which has the ability to perform any operation on the received problem. This paper assumes that adversaries eavesdrop on intercepted communications and pass them forward.
With the growing number of cloud service providers, the threat is increasing. Service providers are in high demand, and their availability is severely limited. This leads to a situation where a new provider offering competitive services can gain popularity rapidly. The internal structure of cloud services remains a mystery, and it is unknown if there is no eavesdropping between the client and the actual computer. Such a new, untrusted provider may be malicious, and communications between the user and the quantum computer may be intercepted.
@@ 5.1 Example parameterization
[…]
@@
[…]
@@ 5.2 An illustrative attack on a scaled-down instance
$x=(\underset{121}{1},\underset{131}{1},\underset{141}{1},\underset{151}{0},\underset{161}{1},\underset{171}{1},\underset{181}{0},\underset{191}{1},\underset{201}{1},\underset{211}{0},\underset{221}{0},\underset{231}{1}).$ (21)